40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 13 03:00:01 2011, Jenne, Update, IOO, WFS shenanigans D990249-B-1_MCWFS_schematic.pdf
    Reply  Thu Jan 13 12:56:57 2011, rana, Update, IOO, WFS shenanigans PerkinElmerQPDs.pdf
    Reply  Fri Jan 28 13:07:31 2011, Jenne, Update, IOO, Beam is back on the WFS 
    Reply  Fri Jan 28 18:19:21 2011, Jenne, Update, IOO, WFS2 has some kind of oil on it Oily_WFS2.jpg
    Reply  Mon Feb 14 15:59:49 2011, Jenne, Update, IOO, WFS quantum efficiency as a function of angle 
       Reply  Wed Jun 15 00:49:34 2011, Suresh, Update, IOO, WFS2 has been fixed. WFS_QE_measurement.pngSensorsBrochure-p12.pdfP6150121.JPGP6150124.JPG
Message ID: 4148     Entry time: Thu Jan 13 03:00:01 2011     Reply to this: 4149   4222   4224   4289
Author: Jenne 
Type: Update 
Category: IOO 
Subject: WFS shenanigans 

My goal this afternoon was to measure the quantum efficiency of the MC WFS.  In the process of doing this, I discovered that when I reverted a change in the MCWFS path (see elog 4107 re: this change), I had not checked the max power going to the WFS when the MC unlocks.

Current status:

MC locks (is locked now).  No light going to WFS at all (to prevent MC WFS french-fry action).  Quantum Efficiency measured.

The Full Story:

Power to WFS:

Rana asked me to check out the quantum efficiency of the WFS, so that we can consider using them for aLIGO.  This involves measuring the power incident on the PDs, and while doing so, I noticed that WFS1 had ~160mW incident and WFS2 had ~240mW incident while the mode cleaner was unlocked.  This is bad, since they should have a max of ~10mW ever.  Not that 200mW is going to destroy the PD immediately, but rather the current out, with the 100V bias that the WFS have, is a truckload of power, and the WFS were in fact getting pretty warm to the touch.  Not so good, if things start melting / failing due to extended exposure to too much heat.

The reason so much power was going to the WFS is that it looks like Yuta/Koji et. al., when trying to use the WFS as a MC1 oplev, changed out 2 of the beam splitters in the MC WFS / MC Refl path, not just one.  Or, we've just been crispy-frying our WFS for a long time.  Who knows?  If it is option A, then it wasn't elogged.  The elog 3878 re: BS changeout only mentions the change of one BS.

Since the MC Refl path has a little more than ~1W of power when the MC is unlocked, and the first BS (which was reverted in elog 4107) is a 10% reflector, so ~100mW goes to the MC Refl PD, and ~900mW goes to the MC WFS path.  In front of a Black Hole beam dump was sitting a BS1-33, so we were getting ~300mW reflected to be split between the 2 WFS, and ~600mW dumped.  The new plan is to put a W2 window in place of this BS1-33, so that we get hopefully something like 0.1% reflected toward the WFS, and everything else will be dumped.  I could not find a W2-45S (everything else is S, so this needs to be S as well).  I found a bunch of W2-0deg, and a few W2-45P.  Does anyone have a secret stash of W2-45S's???  To avoid any more excessive heat just in case, for tonight, I have just left out this mirror entirely, so the whole MC WFS beam is dumped in the Black Hole.  The WFS also have aluminum beam dumps in front of them to prevent light going in.  None of this affects the MC Refl path, so the MC can still lock nice and happily.

Quantum Efficiency Measurement:

I refer to Jamie's LHO elog for the equation governing quantum efficiency of photodiodes: LHO 2 Sept 2009

The information I gathered for each quadrant of each WFS was: [1] Power of light incident on PD (measured with the Ophir power meter), [2] Power of light reflected off the PD (since this light doesn't get absorbed, it's not part of the QE), and [3] the photo current output by the PD (To get this, I measured the voltage out of the DC path that is meant to go to EPICS, and backed out what the current is, based on the schematic, attached). 

I found a nifty 25 pin Dsub breakout board, that you can put in like a cable extension, and you can use clip doodles to look at any of the pins on the cable.  Since this was a PD activity, and I didn't want to die from the 100V bias, I covered all of the pins I wasn't going to use with electrical tape.  After turning down the 100V Kepco that supplies the WFS bias, I stuck the breakout board in the WFS.  Since I was able to measure the voltage at the output of the DC path, if you look at the schematic, I needed to divide this by 2 (to undo the 2nd op amp's gain of 2), and then convert to current using the 499 Ohm resistor, R66 in the 1st DC path.  

I did all 4 quadrants of WFS1 using a 532nm laser pointer, just to make sure that I had my measurement procedure under control, since silicon PDs are nice and sensitive to green.  I got an average QE of ~65% for green, which is not too far off the spec of 70% that Suresh found.

I then did all 8 WFS quadrants using the 1064nm CrystaLaser #2, and got an average QE of ~62% for 1064 (58% if I exclude 2 of the quadrants....see below).  Statistics, and whatever else is needed can wait for tomorrow.

Problem with 2 quadrants of WFS2?

While doing all of this, I noticed that quadrants 3 and 4 of WFS2 seem to be different than all the rest.  You can see this on the MEDM screens in that all 6 other quadrants, when there is no light, read about -0.2, whereas the 2 funny quadrants read positive values.  This might be okay, because they both respond to light, in some kind of proportion to the amount of light on them.  I ended up getting QE of ~72% for both of these quadrants, which doesn't make a whole lot of sense since the spec for green is 70%, and silicon is supposed to be less good for infrared than green.  Anyhow, we'll have to meditate on this.  We should also see if we have a trend, to check how long they have been funny.

Attachment 1: D990249-B-1_MCWFS_schematic.pdf  103 kB  | Show | Show all
ELOG V3.1.3-